# An Industry Interview Study of Software Signing for Supply Chain Security

### v1

@@ 9 Acknowledgments
We thank the study participants for contributing their time to this work. Taylor Schorlemmer and Wenxin Jiang assisted in typesetting. We acknowledge support from Google, Cisco, and NSF #2229703.
Appendix A Outline of Appendices
The appendix contains the following material:
  • Appendix B: The interview protocol.
  • Appendix C: The codebooks used in our analysis, with illustrative quotes mapped to each code.
  • Appendix D: Analysis of saturation in our data.
  • Appendix E: Uncommon challenges, omitted from the main paper due to space limitations.
Appendix B Interview Protocol
@@
Table I gave a summary of the interview protocol. Here we describe the full protocol. We indicate the structured questions (asked of all users) and examples of follow-up questions posed in the semi-structured style. Given the nature of a semi-structured interview, the questions may not be asked exactly as written or in the same sequence but may be adjusted depending on the flow of the conversation.
@@
TABLE VII: Our full interview protocol. The table also indicates the history of the protocol development: removed (strikethrough), changed ($\star$) and re-arranged ($\dagger$) questions.
@@
C: Software Signing in Mitigating Software Supply Chain Risks C-1: If no incident has been recorded (depending on the answer from  LABEL:sec:_B-sec-experienced): why did the team choose to implement software signing? Is software signing the team’s major strategy to secure its supply chain? Any other complimentary security efforts and methods? Are these strategies (software signing and complementary security techniques) influenced by regulations and standards? What challenges or obstacles did your team face while integrating Software Signing into your supply chain security practices? $\star$ How do different team members contribute to the implementation of Software Signing? What roles and responsibilities are involved? (After Practise). Possible Follow up – Do you think Software Signing on its merit (if implemented right) is good enough to completely secure a supply chain? (Added after Pilot).
@@
C-8: $\star$ How does the team evaluate the effectiveness of their Security processes/Signing Implementations? Any Feedback mechanism?
@@
D-6: Have you/your team considered switching this tool for another? AWhat other tools have been considered or are currently being considered?
Appendix C Codebooks
Table VIII describes our codebook, with code, definition, and example quote tagged with that code.
TABLE VIII: Excerpts from the final codebook we used for our analysis. Sample codes under each theme give examples. Note that the theme, “Others and Specific tooling” is not discussed in the paper.
@@ Appendix C Codebooks
Appendix D Saturation Analysis
Figure 5 shows the saturation curve associated with our interviews.
Figure 5: Saturation curve. Each interview covered many topics in detail (orange line, blue line). However, as the green line shows, our results saturated (i.e., stopped observing new perspectives) around interview 11.
Appendix E Uncommon Challenges
E.0.1 Managing Signed Artifacts
E.0.2 Identity Management/Authentication
{quoting}

### v2

@@ 8 Acknowledgments
We thank the study participants for contributing their time to this work. Taylor Schorlemmer, Sofia Okorafor, and Wenxin Jiang assisted in typesetting. We acknowledge support from Google, Cisco, and NSF #2229703.
9 Ethical Considerations
This work considers the human and organizational factors governing the adoption of software signing techniques in software engineering work. To elicit this data, we developed a research protocol, including subject recruitment, interview content, and data processing and analysis. This protocol was approved by our organization’s Institutional Review Board (IRB) prior to the study conduct. All members of the research team who interacted with subjects or their non-anonymized data have completed our organization’s training procedures for handling personally identifiable information. Data from individual subjects were anonymized following best practices for qualitative data analysis.
Any human-subjects research effort must weigh the risks and benefits for individual research subjects as well as for society. The primary risk for subjects is the hazard to their individual or organizational reputation when describing what might be interpreted as sub-standard cybersecurity practices. We therefore anonymized subjects by rank and by organization. A secondary cost for subjects was their time. Subjects were not coerced to participate, and we offered an incentive of \$100 for their participation. The primary potential benefit for our subjects, their organizations, and the societies in which they operate, is that the resulting data may motivate improvements in the software signing ecosystem, as we discussed in section 6. In our judgment, the potential benefit to software cybersecurity far outweighs the risks posed to our subjects and their organizations.
We therefore attest that:
  • The research team considered the ethics of this research, that the authors believe the research was done ethically, and that the team’s next-step plans (e.g., after publication) are ethical.
10 Compliance with Open Science Policy
We acknowledge the commitment to open science, where authors are generally expected to share their research artifacts openly. The research artifacts associated with this study are
  • Raw transcripts of interviews
  • Anonymized transcripts of interviews
  • Interview protocol
  • Codebook
Things we have shared: The interview protocol is a crucial part of any interview study, since it allows for the critical review of a study design as well as its replication. Our full interview protocol is included in Appendix E. Since it is semi-structured, we include all of the questions asked of all subjects, as well as examples of the follow-up questions we asked. We also share the codebook, with codes, definitions, and example quotes that we coded for each code — see Appendix F.
Our full codebook is available at https://doi.org/10.5281/zenodo.14660193.
Things we cannot share: For subject privacy reasons, and for IRB compliance, we cannot share the raw transcripts. We are also unwilling to share the anonymized transcripts of the interviews. Given the high organizational ranks of many of our subjects, and the small size of the subject pool resulting from our recruiting strategy, we believe there is a high risk of de-anonymization even of anonymized transcripts. Therefore, we cannot share the anonymized transcripts.
Appendix A Traditional Software Signing Workflow
We show a typical traditional package signing workflow in Figure 4.
Appendix B Organizational Demographics
We summarize additional organizational demographic information of our subjects in Table 6. This table provides further context for the subject’s organizational size, type of software product, and how many of our subjects belonged to each organization.
Table 6: Organizational Demographics. To enhance anonymity, We only highlight the number of subjects in each organizational category. letters A-L is used to depict each organization.
Type Breakdown (#Organizations|#Subjects)
Organizational Size (Employee Size) Small (<100) (4/6), Medium (<1500) (3/4), Large (>1500) (6/8)
Product Area Digital technology (1/3), SSC Security (2/4), Social technology (1/1), Dev tools (1/1), Telecommunications (1/1), Cloud security (2/3), Aerospace Security (1/1), Internet services (2/2), Cloud + OSS Security (1/1), Cloud + Dev toools (1/1)
Subject Distribution A (3), B (3), C (1), D (1), E (2), F (1), G (1), H (1), I (1), J (1), J (1), K (1), L (1)
Appendix C Code Saturation Curve
We present the trends from the data analysis of our interview subjects in Figure 5. The total number of codes per interview is shown by the blue trend line, with a median of 84.5 codes per interview. Participant S8 had the fewest codes (50), while S7 had the most (130).
The orange trend line represents the unique number of codes in each interview, with a median of 33 unique codes. Subject S4 had the fewest unique codes (24).
The green trend line illustrates the saturation curve for the study, indicating that no new unique codes were observed beyond interview 11 (S11).
As noted earlier (section 4.3), the pilot interviews (S1 and S2) were included due to minimal changes in the interview protocol. These interviews also contained more unique and total codes than the study’s median values, further justifying their inclusion.
Figure 5: Saturation curve. Interviews are plotted in the order in which they were conducted. Each interview covered many topics in detail (orange line, blue line). However, as the green line shows, our results saturated (i.e., stopped observing new perspectives) around interview 11.
Appendix D Other Less Reported Technical Challenges
D.0.1 Managing Signed Artifacts
D.0.2 Identity Management/Authentication
{quoting}
Appendix E Interview Protocol
@@
Table 1 gave a summary of the interview protocol. Here we describe the full protocol. We indicate the structured questions (asked of all users) and examples of follow-up questions posed in the semi-structured style. Given the nature of a semi-structured interview, the questions may not be asked exactly as written or in the same sequence but may be adjusted depending on the flow of the conversation.
@@
Table 7: Our full interview protocol. The table also indicates the history of the protocol development: removed (strikethrough), changed ($\star$) and re-arranged ($\dagger$) questions. We did not analyze part D in this study. We excluded this data due to limited detail and confidentiality concerns. Following qualitative reporting best practices  [85], we disclose this topic to acknowledge its potential influence on data collection.
@@
C: Software Signing in Mitigating Software Supply Chain Risks C-1: If no incident has been recorded (depending on the answer from Section B): why did the team choose to implement software signing? Is software signing the team’s major strategy to secure its supply chain? Any other complimentary security efforts and methods? Are these strategies (software signing and complementary security techniques) influenced by regulations and standards? What challenges or obstacles did your team face while integrating Software Signing into your supply chain security practices? $\star$ How do different team members contribute to the implementation of Software Signing? What roles and responsibilities are involved? (After Practise). Possible Follow up – Do you think Software Signing on its merit (if implemented right) is good enough to completely secure a supply chain? (Added after Pilot).
@@
C-8: $\star$ How does the team evaluate the effectiveness of their Security processes/Signing Implementations? Any Feedback mechanism? (This question is derived from  [39]).
@@
D-6: Have you/your team considered switching this tool for another? What other tools have been considered or are currently being considered?
Appendix F Codebooks
Table 8 describes our codebook, with code definition, and example quote tagged with that code.
Table 8: Excerpts from the final codebook we used for our analysis. Sample codes under each theme give examples. Note that the theme, “Others and Specific tooling” is not discussed in the paper.
@@ Appendix F Codebooks
Appendix G Code Evolution Diagram
Figure 6 describes the evolution of our codebook over five rounds of revisions.
Figure 6: Codebook Evolution: The codebook underwent five rounds of revisions, with changes highlighted by different colors: round 2 (pink), round 3 (blue), round 4 (green), and round 5 (yellow). Arrows indicate the merger or replacement of codes. The final codes from round 5 were used in this paper, and categorized thematically. We combined framework and thematic analyses to develop the codebook. First, we induced codes from each transcript and analyzed them thematically. During the framework analysis, we mapped codes to the four stages/blocks of the software supply chain factory model. Round 1: The codebook was defined based on emerging themes. Round 2: After the intercoder reliability assessment, two categories were merged. Round 3: Blue-highlighted codes (6) were added after memoing and coding the second set of transcripts. Round 4: Green-highlighted codes (4) were added. Round 5: Analysts A1 and A2 determined that all new codes from stages 3 and 4 (except three(3)) were minor and did not fit the established categories. We identified the following themes through framework analysis: Build, Source Code, Deployment, and Third-Party Dependency. Additionally, through thematic analysis, we derived themes such as the Effect of SSC Failures on Signing Adoption, Perceived Importance of Signing, Influence of SSC Standards and Regulations on Signing, and Challenges.